How to find the intercepts for the graph of the equation given. 7x-2y=14

Answer by alicealc(293)

7x-2y=14
x intercept : y = 0
7x = 14
x = 14/7 = 2

y intercept : x = 0
-2y = 14
y= 14/(-2) = -7

so, the x intercept of the graph is (2, 0) and the y intercept of the graph is (0, -7)